OpenStudy (anonymous):

if a single die id rolled find the probability of rolling an even number or a number less than 5

OpenStudy (neer2890):

probability=(no of favorable outcomes)/(no of total outcomes) for even number to come it should be 2,4 or 6(3 cases) and total 6 cases probability=3/6=1/2